# Loop Algorithm for Quantum Transverse Ising Model in a Longitudinal Field

### v1

@@ Abstract
Kagome Qubit Ice.– Constructed with the corner shared triangle, the Kagome lattice exhibits strong geometry frustration so that the local Ising interaction leads to the ground state with macroscopic degeneracy following the ice rule Moessner and Sondhi (2001a). After turning on both transverse and longitudinal fields, the degeneracy will be lifted and the ground state will change to the valence bond solid (VBS) with spontaneous translational symmetry breaking. The presence of the VBS phase is due to the so-called order-by-disorder mechanism and related to the emergent lattice gauge theory with fractional excitations Moessner and Sondhi (2001b); Samajdar et al. (2023); Narasimhan et al. (2024); Lopez-Bezanilla et al. (2023). The QTIM in the Kagome lattice has been realized in the D-WAVE platform where named Kagome qubit ice Narasimhan et al. (2024); Lopez-Bezanilla et al. (2023), and we take it as our final challenging test.
@@
The order parameter to describe the VBS phase is the structure factor defined as $S(\mathrm{Q})=\sum_{i,j}e^{i\mathrm{Q}r_{ij}}\sigma^{(3)}_{i}\sigma^{(3)}_{j}/N$ where $\mathrm{Q}=(\frac{4\pi}{3},0)$. As shown in Fig.4, the structure factors $S(\mathrm{Q})$ calculated by both algorithms match well and reach the maximum value around magnetization per site $M/N=\sum_{i}\sigma^{(3)}_{i}/N$ equals to 1/3, which indicates the formation of the VBS phase. Different from the XXZ model in the Kagome lattice Zhang and Eggert (2013); Zhang et al. (2018); Liu et al. (2024b), the 1/3 magnetization plateau is not flat due to the nonconservation of the magnetization. The ratio of the auto-correlation times in Fig.4(b) strongly supports the advantages of the loop algorithm at the large longitudinal field. However, compared with the Rydberg atom chain, such advantages become not obvious in the small longitudinal field and may be due to the high degeneracy of the ground and excited quantum states.
@@
Figure 4: Kagome qubit ice. (a) Structure factor and (b) the ratio of its auto-correlation time by line and loop algorithm. The inset of (a) shows the magnetization per site and the dashed black line highlights the 1/3 corresponding to the ice rule filling. The parameters are $J_{z}=0.25$, $\Gamma=0.15$, $\beta=20$, and $N=24\times 24\times 3$ with periodical boundary condition.
@@
Conclusion and discussion– By innovatively introducing a novel update strategy: merge-unmerge process, We successfully design a loop-type QMC algorithm tailored for the QTIM and its extensions. After comparing with the state-of-the-art line cluster method by simulating two typical realistic platforms, the loop algorithm demonstrates significant advantages, especially in the large longitudinal field region. Meanwhile, the loop algorithm is more coding-friendly. The loop algorithm cannot only be taken as advantageous numerical simulators for quantum magnetism, trapped ion, Rydberg array, quantum computer, and many other quantum many-body systems but also easily extended to more complex models in statistic physics, like the quantum clock and Potts models. Furthermore, such merge-unmerge update process can also be ported to the continuous time worm algorithm Liu et al. (2020b).
@@
Acknowledgement– X.-F. Z. thanks Zheng Yan for fruitful discussions and acknowledges funding from the National Science Foundation of China under Grants No. 12274046, No. 11874094, No.12147102 and No.12347101, Chongqing Natural Science Foundation under Grants No. CSTB2022NSCQ-JQX0018, Fundamental Research Funds for the Central Universities Grant No. 2021CDJZYJH-003, and Xiaomi Foundation / Xiaomi Young Talents Program.

### v2

@@ III.2 Kagome Qubit Ice
Constructed with the corner shared triangle, the Kagome lattice exhibits strong geometry frustration so that the local Ising interaction leads to the ground state with macroscopic degeneracy following the ice rule Moessner and Sondhi (2001a). After turning on both transverse and longitudinal fields, the degeneracy will be lifted and the ground state will change to the valence bond solid (VBS) with spontaneous translational symmetry breaking. The presence of the VBS phase is due to the so-called order-by-disorder mechanism and related to the emergent lattice gauge theory with fractional excitations Moessner and Sondhi (2001b); Samajdar et al. (2023); Narasimhan et al. (2024); Lopez-Bezanilla et al. (2023). The QTIM in the Kagome lattice has been realized in the D-WAVE platform where named Kagome qubit ice Narasimhan et al. (2024); Lopez-Bezanilla et al. (2023), and we take it as our final challenging test.
@@
The order parameter to describe the VBS phase is the structure factor defined as $S(\mathrm{Q})=\sum_{i,j}e^{i\mathrm{Q}r_{ij}}\sigma^{(3)}_{i}\sigma^{(3)}_{j}/N$ where $\mathrm{Q}=(\frac{4\pi}{3},0)$. As shown in Fig. 4, the structure factors $S(\mathrm{Q})$ calculated by both algorithms match well and reach the maximum value around magnetization per site $M/N=\sum_{i}\sigma^{(3)}_{i}/N$ equals to 1/3, which indicates the formation of the VBS phase. Different from the XXZ model in the Kagome lattice Zhang and Eggert (2013); Zhang et al. (2018); Liu et al. (2024b), the 1/3 magnetization plateau is not flat due to the nonconservation of the magnetization. The ratio of the auto-correlation times in Fig. 4(b) strongly supports the advantages of the loop algorithm at the large longitudinal field. Consistent with observations in Rydberg atom arrays, the loop algorithm demonstrates superior performance compared to both the line algorithm and conventional loop algorithms across the entire parameter range investigated. Specifically, for $B>0.5$, the auto-correlation time of the loop algorithm is 2-3 times shorter than that of the line algorithm. When compared to conventional loop algorithms, the reduction in auto-correlation time reaches up to two orders of magnitude, with this advantage becoming increasingly pronounced at higher longitudinal field strengths. However, compared with the Rydberg atom chain, such advantages become not obvious in the small longitudinal field and may be due to the high degeneracy of the ground and excited quantum states. While introducing the second and third nearest neighbor interactions, our recent work demonstrates that geometric string breaking can be observed by tailoring the open edge in the Kagome Rydberg arrays Xu and Zhang (2025).
@@
Figure 4: Kagome qubit ice. (a) Structure factor and (b) the ratio of its auto-correlation time by line and loop algorithm. The inset of (a) shows the magnetization per site and the dashed black line highlights the 1/3 corresponding to the ice rule filling. The inset of (b) shows the ratio of its auto-correlation time by $\mathrm{loop_{c}}$ and loop algorithm. The parameters are $J_{z}=0.25$, $\Gamma=0.15$, $\beta=20$, and $N=24\times 24\times 3$ with periodical boundary condition.
IV Conclusion and discussion
@@
By innovatively inventing a novel update strategy: merge-unmerge process, we successfully design a loop-type QMC algorithm tailored for the QTIM and its extensions. The merge-unmerge process can cause the loop update path to move in the spatial direction in the extended configuration space, while also avoiding the decrease in operator weight that occurs due to subdivision in the directed loop algorithm. After comparing with the state-of-the-art line cluster method by simulating two typical realistic platforms, the loop algorithm demonstrates significant advantages, especially in the large longitudinal field region. Meanwhile, the loop algorithm is more coding-friendly and the teaching code is open-source sup .
The loop algorithm can be viewed not only as an advanced numerical simulator for quantum magnetism, trapped ions, Rydberg atom arrays, quantum computers, and many other quantum many-body systems, but also easily extended to more complex models in statistical physics, such as the quantum clock and Potts models. Furthermore, such a merge-unmerge update process can also be ported to other Monte-Carlo methods, such as the continuous-time worm algorithm Liu et al. (2020b). Meanwhile, many advanced techniques related to the entanglement measurement Wang et al. (2025); Ding et al. (2025a, b); Mao et al. (2025); Ding et al. (2024) can also be incorporated into our algorithm.
V Acknowledgment
@@
X.-F. Z. thanks Zheng Yan for fruitful discussions, MPI-PKS for hospitality during the visit, and acknowledges funding from the National Science Foundation of China under Grants No. 12274046, No. 11874094, No.12147102, and No.12347101, Chongqing Natural Science Foundation under Grants No. CSTB2022NSCQ-JQX0018, Fundamental Research Funds for the Central Universities Grant No. 2021CDJZYJH-003, and Xiaomi Foundation / Xiaomi Young Talents Program.
Appendix A Detailed Balance Condition
The detailed balance condition of the diagonal update is the same as Ref. Lee et al. (1984). For the loop to pass through the operators, the updating methods are conventional, so the detailed balanced condition is also kept. Therefore, the merge-unmerge process and the starting and stopping of the loop processes should be discussed in detail.
The merge-unmerge process is similar to the link-cluster process in the Swendsen-Wang algorithm and the start and stop processes of the loop update in the directed loop algorithm. The balance condition should be satisfied between the original configuration space and the extended configuration space including merged operators. Because the unmerge process is the exact reverse process of the merge process, the detailed balance condition is kept.
In the loop updating process, the detailed balance can be verified by analyzing a loop update path with its inverse path. Because the number of the merged operators $N_{o}$ is not changed, the probability $1/N_{o}$ due to random selection of the starting position of the loop is equal for all the loops in each loop update. Therefore, it is not necessary to include it into the acceptability of starting or stopping the loop. Then, the possibility that the loop stops after passing the $N_{s}$ merged operator is $(1-P_{s})^{N_{s}}P_{s}$. Because the loop and its reverse process pass the same number of merged operators, this possibility related to the stopping process will not affect the detailed balanced conditions.
Figure 5: The loop length of (a,b) the Rydberg atom chain and (c,d) Kagome qubit ice. (a) The Rabi frequency $\Omega$ is changing at $\Delta=0$ and (b) the detuning $\Delta$ is changing at $\Omega=1.0$. The other parameters are identical to those in Fig. 3: $R_{b}=1.2$, $\beta=20$, and $L=51$. (c) The transverse field $\Gamma$ is changing at BB=0.5 and (d) the longitudinal field $B$ is changing at $\Gamma$=0.15. The other parameters are identical to those in Fig. 4: $J_{z}=0.25$, $\beta=20$ and $N=6\times 6\times 3$.
Appendix B Loop Length Analysis
The loop lengths in both the Rydberg chain and Kagome qubit ice are plotted in Fig. 5. Different from the closed-loop algorithm, the loop length becomes larger with the decrease of the transverse field. That’s because the number of off-diagonal site operators increases while enlarging the transverse field so that the loop meets the merged operator more frequently and stops earlier. Similarly, when increasing the magnitude of the longitudinal field (detuning) with a fixed transverse field (Rabi frequency), the number of diagonal operators strongly increases, consequently, it would be harder for the loop to stop, and the average length of the loop will increase.
However, in Fig. 5(b), we can find the decline of the loop when approaching $\Delta=-3$ (corresponding large negative longitudinal field). The possible explanation is that most of the Rydberg atoms stay in the ground state so that the van der Waals interactions have less effect on the diagonal energy.
In conclusion, different from the loop length in the directed loop algorithm, which is highly relevant to the type of phase (due to its connection to the off-diagonal correlation function), the loop length in our algorithm depends exclusively on the ratio between merged and diagonal operators. That may be the key reason our algorithm shows advantages, especially for a large longitudinal field parameter region.
Figure 6: Comparison of the autocorrelation time $\tau^{\text{loop}}_{\text{mc}}$ versus the stop probability $P_{s}$ at two points of Fig. 3: (a) near the critical point ($\Delta=1$) and (b) deep in the ordered phase ($\Delta=3$). The number of Monte Carlo steps for (b) was increased to $4\times 10^{7}$ to reduce statistical errors.
Appendix C Stop Probability
We investigated the relationship between the stop probability, $P_{s}$, and the autocorrelation time, $\tau^{\text{loop}}_{\text{mc}}$. The typical behavior of this relationship is shown in Fig. 6. To measure it, we selected two parameter points ($\Delta=1$ and $\Delta=3$) from the phase diagram in Fig. 3. All other simulation parameters were identical to those used in Fig. 6, except that the number of Monte Carlo steps for the data in panel (b) was increased to $4\times 10^{7}$ to reduce the statistical error. The parameters corresponding to panel (a) are near the phase transition point, while those for panel (b) lie within the antiferromagnetic phase.
As shown in Fig. 6, $\tau^{\text{loop}}_{\text{mc}}$ initially decreases as $P_{s}$ increases, reaches a minimum around $P_{s}\approx 0.4$, and then increases again for $P_{s}\gtrsim 0.8$. The unbiased stop probability of $P_{s}=0.5$, chosen for our simulations, sits within the region of minimal autocorrelation time.
Our analysis confirms a non-monotonic dependence of the autocorrelation time on the stop probability $P_{s}$, with an optimum near $P_{s}\approx 0.6$. As this optimum is system-dependent, we recommend testing a narrow range of $P_{s}$ values for performance-critical applications.